# A Community Palm Model

### v1

@@ 2.5 Risk assessment
Here we define risk as the sum of potential land cover transitions times their associated probabilities in a region of interest (ROI). To-palm is the transition of non-palm land cover to palm in any pixel, in units of pixel area. From-palm is the transition from palm to any non-palm land cover in any pixel, in units of pixel area. We categorize land cover into two categories: forest and non-forest, according to (Bourgoin et al., 2023). We define the probability of the to-palm transition as P(Pt-1=0, Pt=1) where P is the probability operator, and Pt (italicized) is a Bernoulli distributed random variable that indicates palm presence (1) or absence (0) at time t. The probability of the to-palm transition is the probability of not palm at time t-1 and palm at time t.
The pair of Bernoullis at two points in time (Pt-1, Pt) is a bivariate Bernoulli distributed random variable (Marshall and Olkin, 1985; DAI et al., 2013). Denote the probability of each possible state according to the contingency Table 1:
@@
Assume that the Spearman correlation $(F(x_{t-1}),F(x_{t}))$ in a spatio-temporal neighborhood (i.e. the correlation computed in a 210x210 meter, 21x21 pixel, neighborhood centered on each pixel in the image inputs) is a fair approximation to the Pearson correlation of the unknown indicator variables. Note that the Pearson correlation is defined as:
@@
We define the cost of a transition as the area (hectares) of a pixel $(A_{i,j})$ in which a transition occurred. The risk of transition associated with a given ROI is therefore the sum over all $i,j$ in the ROI of Ai,j*P(transition)i,j.
@@ 3 Results
Figures 3-5 show 2020 annual composites of the predictors in the model at a specific location.
@@
The model is published as a trained TensorFlow model on GitHub (https://github.com/google/forest-data-partnership). For interactive display and/or further analysis, the model can be hosted on platforms like Google Vertex AI and model output accessed from Google Earth Engine. For palm producing countries, annual probability maps for 2020 and 2023 are produced as 10m raster image collections that are available in the Google Earth Engine public data catalog under a CC-BY 4.0 license (see Section 6). Figures 6-7 shows the model output (estimated conditional probabilities) for years 2020 and 2023.
@@ 3.1 Accuracy
Accuracy measures computed from the geographic cross-validation are shown in Table 2. The statistics are roughly stable across folds. All show relatively high precision and lower recall, especially at higher thresholds on the output. Binary accuracy (threshold = 0.5) is approximately 90% on the holdout data.
Table 2: Accuracy over three geographic folds.
Accuracy Measure Fold 0 Fold 1 Fold 2
Cross-entropy loss 0.190 0.213 0.193
Binary accuracy @ 0.5 threshold 0.910 0.904 0.892
Area under the ROC curve (AUC) 0.985 0.976 0.982
Recall @ 0.95 precision 0.960 0.938 0.943
Recall @ 0.5 threshold 0.958 0.932 0.960
Recall @ 0.7 threshold 0.918 0.898 0.924
Recall @ 0.9 threshold 0.805 0.793 0.809
Precision @ 0.5 threshold 0.953 0.954 0.934
Precision @ 0.7 threshold 0.975 0.971 0.962
Precision @ 0.9 threshold 0.990 0.985 0.986
Precision @ 0.95 recall 0.960 0.939 0.944
Fold 0 held out corresponds to the deployed model (see Section 6), which was trained on folds 1 and 2. AUC is approximately 0.985 in this configuration, illustrating relatively strong predictive performance, and is consistently high in the other folds as well. Figure 8 shows the geographic distribution of error in a regional area. Although the error is relatively well-distributed spatially, clustered errors suggest locations where additional training/validation data could improve model performance and characterization.
Figure 8:
The cross-validation illustrates potential self-evaluation bias and other issues. Table 3 shows the performance of the predictor (binary accuracy @ 0.5 threshold) when one entire dataset is withheld. These accuracy measures are lower than the geographic cross-fold validation. The geographic partition indicates binary accuracy (0.5 threshold) at around 90%. Leaving one dataset out, that level of accuracy is only obtained on the Vollrath and Descals datasets. When held out, other datasets show 73-88% accuracy. The model accuracy decreases most significantly when the Google dataset is withheld. The training set accuracy of the withheld datasets is not consistently higher than test set accuracy, as would be expected since the data used to train the model is in the same geographic partition.
Table 3: Accuracy over dataset folds.
Dataset Test Fold Accuracy Train Fold Accuracy
Google 0.747 0.737
Vollrath et al. 0.909 0.898
Danylo et al. 0.797 0.794
Descals et al. 0.928 0.926
T. Lips 0.881 0.779
@@ 3.2 Risk assessment
Figure 9 shows the estimated probability of palm transitions from years 2020 to 2023, computed according to equation 3. The three regions shown in Fig. 9 correspond to hypothetical sourcing domains. Manual inspection of the imagery suggests that the red region represents an area without palm, the green region represents an established palm plantation, and the blue region represents a mixed area. The risk of palm transition associated with each ROI was computed as the sum of estimated probability of the transition in each pixel times cost of the transition (pixel area in hectares). The results are shown in Table 4.
Figure 9: Hypothetical sourcing domains and p01 transition probability.
@@
Forest 290.3 0 80.5
To-palm risk (forest) 27.1 N/A 12.4
From-palm risk (forest) 33.8 N/A 11.4
Non-forest 4.6 325.5 336.1
To-palm risk (non-forest) 0.4 1.5 78.9
From-palm risk (non-forest) 0.6 6.0 28.1
@@ 4 Discussion
Interpreting the model output as probabilities retains local uncertainty in the underlying condition, which is unknown. The global oil palm probability map developed with this method can be a two-class (palm versus not-palm) classification by thresholding probabilities using automatic methods such as Otsu segmentation (Otsu, 1979). The threshold determined by Otsu for the 2023 palm map (roughly the area of Figures 6-7) is 0.515. From Table 2, both precision and recall are relatively high ( 95%) at a 0.5 threshold. Thresholding the 2023 palm in Figure 7 at 0.515 indicates 5095 hectares of planted oil palm in the area of Figure 7. Any selected threshold represents a trade-off between recall and precision as shown in Table 2. For this reason, it is incumbent on the user of the data to determine a threshold suitable for their use case and region of interest, for example using Otsu segmentation to select the best threshold for local conditions.
However, threshold delineation is not required, as the probabilities can be used directly. Treating the probabilities as an estimated distribution over the unknown (and assumed hidden) Bernoulli random variable (Palm: 0, 1), indicates 5008 hectares of planted palm in Figure 7 – a 1.7% difference compared to the threshold method. Since the contingency table can be specified completely with the assumptions described previously, it can be used to generate estimates of risk. These risk estimates are probabilities and account for the uncertainty in model estimates at multiple points in time. The risks can be used directly in supply chain analysis to identify areas of palm transitions that indicate the need for additional verification or characterization.
@@
For example, consider the three regions in Figure 9 and Table 4. In region 1, the majority of the area is in forest, as defined by Bourgoin et al. (2023). Although the to-palm risk is  27 hectares, there is also from-palm risk of  34 hectares. In region 1, we conclude that the model output is noisy, as the from-palm transition risk is larger than the to-palm transition risk. In region 2, there is no forest, and small amounts of transition risk, which we conclude to be either noise or active palm plantation management. In region 3, we note substantial areas of both forest and non-forest. In the non-forest area of region 3, there are relatively high amounts of transition, which we assume to be palm management. In the forested area of region 3, there is  12 hectares of to-palm risk and  11 hectares of from-palm risk. For this reason, we conclude that the risk of palm conversion is highest in region 3, with at least one hectare of risk. It is incumbent upon the user to decide where such transitions are problematic, for example by incorporation of a forest layer. Limitations of this approach are discussed below.
@@
For such a model to be effective in a regulatory context, oil palm maps must be produced on a yearly basis, since compliance is determined annually. The model also needs to be easily updatable, meaning that stakeholders can contribute data in their areas of interest, if the model output is not representative of that area. This is especially important from an ethical perspective, where local stakeholders need inclusion in a process that could affect their livelihoods. Further verification of submissions may be required where submissions are found to reduce overall or local model accuracy. A single iteration of the process consists of an examination of model results by a relevant stakeholder, creation of additional training/validation data in the stakeholder’s ROI, model refinement and retraining using the new data, and publication of the updated model. That workflow was demonstrated here. Specifically, preliminary model results were provided to T. Lips, who created a new training dataset, after which point the improved model was retrained and republished. Additional iterations and version tracking are needed to keep the model inclusive, accurate and up-to-date.
@@ 4.1 Limitations
The model output quality could be limited by the expressive power of the neural network itself, or the specific choices of inputs. Although uncertainty is implicit in the model output, which is itself expressed as a probability, the model could be biased due to training data biases or an inability to generalize the training data in the model weights. This bias could be reduced through enhancements to model architecture or its set of input features. For example, both could be enhanced through inputting spatio-temporal patches of inputs instead of single pixel vectors from annual composites. There could also be high variance in the model output. This is illustrated by the analysis of region 1, in which noise in model output is assumed to account for both from- and to-palm transition in the region. Variance could be reduced through collection of a more geographically representative and balanced training data set. Due to known limitations resulting from model bias and/or variance, quality of predictions will vary geographically and results should be interpreted with caution. We expect more training data, model tuning, and/or improved model architecture to improve model accuracy. For optimization of model performance and additional data collection, diagnosing the balance of bias and variance in model error will be the subject of future research.
The probabilistic framework on which the risk assessment is based relies on strong assumptions. We assume that the true state of a pixel can be adequately represented by a Bernoulli random variable, implying that a pixel can be in only one of two states: palm or not palm. We make little effort to determine, or even to define, such states, which are simply taught to the model directly from the data. Given a vector of predictors, we assume the model output is a suitable representation of the conditional probability of oil palm at a given location, and treat it as the marginal probability of the unknown indicator variable. Finally, the estimate of palm conversion depends on the estimation of correlation of the unknown indicators from the Pearson rank correlation of the model outputs. The Pearson correlation is probably an overestimate of the Spearman correlation between the unknown indicators, since the rank-correlation can be high for non-linearly related variables. This would overestimate p11 and consequently underestimate the transition risk.
@@ 4.2 Future Work
We intend for the model presented here to continuously improve. To maximize effectiveness, the model should evolve as relevant community stakeholders present it with additional data. To this end, we invite the larger community of local, regional, and global stakeholders to submit informed feedback, additional training data and/or other information (feedback form: https://forms.gle/KzGgGPn4pLYydJdHA, or the Collect Earth Online data collection tool https://app.collect.earth/review-institution?institutionId=2494). Improvements are not limited to a steadily increasing amount of training and/or validation data. The model could be improved through changes in input format (e.g. patches, which contain more spatial information) and/or model architecture. The inputs could be improved through incorporation of more information as new sensors come online or through learned embedding spaces ((Jean et al., 2019)). Community contributions to model and or input data are welcome through our GitHub repository (see Section 6). We expect modeling other land covers of interest (e.g. cocoa, coffee, rubber) in a community framework to increase accuracy and availability of information about the spatio-temporal distribution of commodities and relationship to deforestation.
@@ 5 Conclusion
This model, and resulting oil palm probability maps are useful for accurately identifying the geographic footprint of oil palm cultivation, and how that footprint is changing over time. We have demonstrated how the analysis of the model outputs at multiple points in time can be used to infer oil palm dynamics. We recognize that the commodity probability maps are not solely sufficient to quantify palm oil driven deforestation, however they are necessary to disambiguate deforestation of natural forests from other kinds of land cover transitions. Used in conjunction with other geospatial information, such as global forest cover, deforestation alerts and high resolution imagery, this community palm oil map and risk indicators will support organizations’ understandings of the impact of their supply chains, help to verify information provided by producers, and support due diligence reporting. Specifically for regulatory mechanisms like the EUDR, we intend for the community model and resulting data products support the “convergence of evidence” approach developed by the Forest Data Partnership to support due diligence statements and compliance claims.
@@ 6 Model and Data
Earth Engine assets:
ee.ImageCollection("projects/forestdatapartnership/assets/palm/palm_2020_model_20240312")
ee.ImageCollection("projects/forestdatapartnership/assets/palm/palm_2023_model_20240312")

### v2

@@ 2.5 Risk assessment
Here we define risk as the sum of potential land cover transitions times their associated probabilities in a region of interest (ROI). To-palm is the transition of non-palm land cover to palm in any pixel, in units of pixel area. From-palm is the transition from palm to any non-palm land cover in any pixel, in units of pixel area. We categorize land cover into two categories: forest and non-forest, according to Bourgoin et al. (2023). We define the probability of the to-palm transition as $\textrm{P}(P_{t-1}=0,P_{t}=1)$ where P is the probability operator, and $P_{t}$ (italicized) is a Bernoulli distributed random variable that indicates palm presence (1) or absence (0) at time $t$. The probability of the to-palm transition is the probability of not palm at time $t-1$ and palm at time $t$.
The pair of Bernoullis at two points in time $(P_{t-1},P_{t})$ is a bivariate Bernoulli distributed random variable (Marshall and Olkin, 1985; DAI et al., 2013). Denote the probability of each possible state according to the contingency Table 1:
@@
Assume that the Spearman correlation $(F(x_{t-1}),F(x_{t}))$ in a spatio-temporal neighborhood (i.e. the correlation computed in a 310x310 meter, 31x31 pixel, neighborhood centered on each pixel in the image inputs) is a fair approximation to the Pearson correlation of the unknown indicator variables. Note that the Pearson correlation is defined as:
@@
Note that we can also estimate p10 (P(from-palm)) by substituting $P_{t-1}$ for $P_{t}$ in equation 3. We define the cost of a transition as the area (hectares) of a pixel $(A_{i,j})$ in which a transition occurred. The risk of transition associated with a given ROI is therefore the sum over all $i,j$ in the ROI of Ai,j*P(transition)i,j.
@@ 3 Results
Figures 3-6 show 2020 annual composites of the predictors in the model at a specific location.
@@
Figure 6: Annual (2020) Palsar-2 composite, HH, HV, HH bands as red, green blue.
The model is published as a trained TensorFlow model on GitHub (https://github.com/google/forest-data-partnership). For interactive display and/or further analysis, the model can be hosted on platforms like Google Vertex AI and model output accessed from Google Earth Engine.
For palm producing countries, annual probability maps for 2020 and 2023 are produced as 10m raster image collections that are available in the Google Earth Engine public data catalog under a CC-BY 4.0 NC license for non-commercial users and commercial licensing by request (see Section 6). Figures 6-7 shows the model output (estimated conditional probabilities) for years 2020 and 2023.
@@ 3.1 Accuracy
Table 2 shows accuracy measures computed from the geographic test fold and folds 0 and 2 of the Descals et al. (2024) validation set. Binary accuracy (threshold = 0.5) is approximately 92% on the holdout data.
Table 2: Accuracy measures on multiple datasets.
Accuracy Measure Test fold Descals et al. 2024
Cross-entropy loss 0.097 0.183
Binary accuracy @ 0.5 threshold 0.969 0.924
Area under the ROC curve (AUC) 0.992 0.976
Recall @ 0.5 threshold 0.968 0.913
Precision @ 0.5 threshold 0.969 0.600
The accuracy assessment illustrates several important issues. The test fold consistently shows higher accuracy than the Descals et al. (2024) validation set, suggesting potential self-evaluation bias and/or dependence between folds. Since the Descals et al. (2024) validation data were not seen in model training or selection, and were restricted to geographic locations different from the training data, these are a more conservative estimate of overall model performance. Note that the precision at threshold 0.5 is  60% on this set, suggesting overestimation in non-palm areas at low thresholds. Figure 9 shows overall accuracy, F1, precision and recall curves for the model on the Descals et al. (2024) validation set. Note that the threshold which optimizes F1 score is closer to 0.82, where precision and recall are  82%. Figure 9 also illustrates the extreme imbalance (towards negatives) of the the Descals et al. (2024) validation set. Specifically, overall accuracy increases with increasing threshold, which is near 96% at the 0.82 threshold.
Figure 9: Accuracy, F1, precision and recall curves on Descals et al. 2024 validation data.
@@ 3.2 Risk assessment
Figure 10 shows the estimated probability of palm transitions from years 2020 to 2023, computed according to equation 3 or its analog for $P_{t-1}$. The three regions shown in Fig. 10 correspond to hypothetical sourcing domains. Inspection of the imagery suggests that region 1 (red) represents an area without palm, the region 2 (green) represents a mixed area, and region 3 (blue) represents an established palm plantation. The risk of palm transition associated with each ROI was computed as the sum of estimated probability of the transition in each pixel times cost of the transition (pixel area in hectares). The results are shown in Table 3.
Figure 10: Hypothetical sourcing domains and palm transition probabilities, with 2023 Sentinel-2 RGB background.
@@
Forest 983 102 0
To-palm risk (forest) 17.9 2.4 N/A
From-palm risk (forest) 14.9 1.7 N/A
Non-forest 331 1465 1266
To-palm risk (non-forest) 5.1 83.6 4.8
From-palm risk (non-forest) 5.3 126.5 32.9
@@ 4 Discussion
Interpreting the model output as probabilities retains local uncertainty in the underlying condition, which is unknown. The global oil palm probability map developed with this method can be a two-class (palm versus not-palm) classification by thresholding probabilities using automatic methods such as Otsu segmentation (Otsu, 1979). The threshold determined by Otsu for the 2023 palm map (roughly the area of Figure 10) is 0.506. Thresholding the 2023 palm in Figure 7 at 0.506 indicates 5842 hectares of planted palm in the area of Figure 10. From Table 2, recall is relatively high at a 0.5 threshold, but precision is low]. Any selected threshold represents a trade-off between recall and precision as shown in Figure 9. For this reason, it is incumbent on the user of the data to determine a threshold suitable for their use case and region of interest, for example using Otsu segmentation to select an optimal threshold for local conditions. However, threshold delineation is not required, as the probabilities can be used directly. Treating the probabilities as an estimated distribution over the unknown (and assumed hidden) Bernoulli random variable (Palm: 0, 1), indicates 6317 hectares of planted palm in Figure 7 – an  8% difference compared to the threshold method.
Figure 11: Estimated palm probability: both 2020 and 2023.
Since the contingency table can be specified completely with the assumptions described previously, it can be used to generate estimates of the probability of different scenarios while accounting for uncertainty in model estimates at multiple points in time. For example, the estimate of p11 is the probability of "stable" palm, or places that were palm in both 2020 and 2023 (Figure 11). Estimates of persistent palm seem less noisy than single year output, but additional testing is required to confirm that. Risk can be computed for any scenario, in units of area, using the method described. Although risks computed in this manner can be used directly in supply chain analysis, due to potential model error, the risks should be interpreted with caution. Specifically, these data can indicate the need for additional verification or characterization of high risk areas, but should be combined with other evidence to establish risk definitively.
@@
For example, consider the three regions in Figure 10 and Table 3. In region 1, the majority of the area is in forest, as defined by Bourgoin et al. (2023). Although the to-palm risk is  18 hectares, there is also from-palm risk of  15 hectares. In region 1, we conclude that the model output is noisy, as the from-palm transition risk roughly balances the to-palm transition risk. In region 2, there is some forest, and small amounts of transition risk, which we conclude to be either noise or active palm plantation management. Region 3 is all non-forest. In region 3, there are relatively high amounts of transition, which we assume to be palm management. We don’t interpret any of these areas as particularly high risk. Limitations of this approach are discussed below.
@@
For such a model to be effective in meeting sustainability commitments, oil palm maps must be produced on a yearly basis, since compliance is determined annually. The model also needs to be easy to update, meaning that stakeholders can contribute data in their areas of interest, if the model output is not representative of that area. This is especially important from an ethical perspective, where local stakeholders need inclusion in a process that could affect their livelihoods. Further verification of submissions may be required where submissions are found to reduce overall or local model accuracy. A single iteration of the process consists of an examination of model results by a relevant stakeholder, creation of additional training/validation data in the stakeholder’s ROI, model refinement and retraining using the new data, and publication of the updated model. That workflow was demonstrated here. Specifically, preliminary model results were provided to T. Lips, who created a new training dataset, after which point the improved model was retrained and republished. Additional iterations and version tracking are needed to keep the model inclusive, accurate and up-to-date.
@@ 4.1 Limitations
Although uncertainty is implicit in the model output, which is expressed as a probability, the model could be biased due to training data biases or an inability to generalize the training data in the model weights. This bias could be reduced through enhancements to model architecture, for example through inputting spatio-temporal patches of inputs instead of single pixel vectors from annual composites. There could also be high variance in the model output. This is illustrated by the analysis of region 1, in which noise in model output is assumed to account for a substantial amount of from- and to-palm transitions in the region. Variance could be reduced through collection of a more geographically representative and balanced training data set. We suspect variance is also reduced by combining multiple years of predictions to isolate stable crops of interest (e.g. Figure 11), but additional research is needed to confirm that suspicion. Presently, known limitations resulting from model bias and/or variance mean quality of predictions will vary geographically and results should be interpreted with caution.
The probabilistic framework on which the risk assessment is based relies on strong assumptions. We assume that the true state of a pixel can be adequately represented by a Bernoulli random variable, implying that a pixel can be in only one of two states: palm or not palm. We make little effort to determine, or even to define, such states, which are simply taught to the model directly from the data. Given a vector of predictors, we assume the model output is a suitable representation of the conditional probability of oil palm at a given location, and treat it as the marginal probability of the unknown indicator variable. Finally, the estimate of palm conversion depends on the estimation of correlation of the unknown indicators from the Spearman rank correlation of the model outputs. The Spearman correlation is probably an overestimate of the Pearson correlation between the unknown indicators, since the rank-correlation can be high for non-linearly related variables. This would overestimate p11 and consequently underestimate the transition risk.
@@ 4.2 Future Work
We intend for the model presented here to continuously improve. To maximize effectiveness, the model should evolve as relevant community stakeholders contribute additional data. To this end, we invite the larger community of local, regional, and global stakeholders to submit informed feedback, additional training data and/or other information (feedback form: https://goo.gle/fdap-data. Improvements are not limited to a steadily increasing amount of training and/or validation data. The model could be improved through changes in input format (e.g. patches, which contain more spatial information) and/or model architecture. The inputs could be improved through incorporation of more information as new sensors come online or through learned embedding spaces ((Jean et al., 2019)). Community contributions to model and or input data are welcome through our GitHub repository (see Section 6). We expect modeling other land covers of interest (e.g. cocoa, coffee, rubber) in a community framework to increase accuracy and availability of information about the spatio-temporal distribution of commodities and relationship to deforestation.
@@ 5 Conclusion
This model, and resulting oil palm probability maps are useful for accurately identifying the geographic footprint of oil palm cultivation, and how that footprint is changing over time. We have demonstrated how the analysis of the model outputs at multiple points in time can be used to infer oil palm dynamics. We recognize that the commodity probability maps are not solely sufficient to quantify palm oil driven deforestation, however they are necessary to disambiguate deforestation of natural forests from other kinds of land cover transitions. Used in conjunction with other geospatial information, such as global forest cover, deforestation alerts and high resolution imagery, this community palm oil map and risk indicators will support organizations’ understanding of the impact of their supply chains, help to verify information provided by producers, and support due diligence reporting. Specifically for regulatory mechanisms like the EUDR, we intend for the community model and resulting data products to support the “convergence of evidence” approach developed by the Forest Data Partnership for due diligence statements and compliance claims.
@@ 6 Model and Data
Earth Engine asset:
ee.ImageCollection("projects/forestdatapartnership/assets/palm/model_2024a")